OpenStudy (anonymous):

If a ball is thrown directly upward with a velocity of 40 ft/s, its heigh in feet after t seconds is given by y=40t-16t^2. What is the maximum height attained by the ball? I am completely lost on this problem.

OpenStudy (anonymous):

\[f(x)=y =ax^2+bx+c\]is the general form for a quadratic. it has a maximum if a<0 and a minimum if a>0. the max/min occurs when \[x=\frac{ -b }{ 2a }\] and the max/min value is \[f \left( \frac{ -b }{ 2a } \right)\] for your problem, \[f(x) = y = -16t^2 +40t\] a = -16 and b = 40 since a < 0 you will have a maximum. it occurs when \[x = \frac{ -b }{ 2a }=\frac{ -(40) }{ 2\cdot (-16) }=\frac{ -40 }{ -32 }=\frac{ 5 }{ 4 }\] so plug in... \[f \left( \frac{ 5 }{ 4 } \right)=-16\left( \frac{ 5 }{ 4 } \right)^2 + 40\left( \frac{ 5 }{ 4 } \right)=-25+50=25\]

OpenStudy (anonymous):

sorry, used x when it should have been t.

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes it does! Thanks!! :)